# Real-Time Recurrent Reinforcement Learning

### v2

@@ Abstract
In this paper we propose real-time recurrent reinforcement learning (RTRRL), a biologically plausible approach to solving discrete and continuous control tasks in partially-observable markov decision processes (POMDPs). RTRRL consists of three parts: (1) a Meta-RL RNN architecture, implementing on its own an actor-critic algorithm; (2) an outer reinforcement learning algorithm, exploiting temporal difference learning and dutch eligibility traces to train the Meta-RL network; and (3) random-feedback local-online (RFLO) learning, an online automatic differentiation algorithm for computing the gradients with respect to parameters of the network. Our experimental results show that by replacing the optimization algorithm in RTRRL with the biologically implausible back propagation through time (BPTT), or real-time recurrent learning (RTRL), one does not improve returns, while matching the computational complexity for BPTT, and even increasing complexity for RTRL. RTRRL thus serves as a model of learning in biological neural networks, mimicking reward pathways in the basal ganglia.
@@ Introduction
Model-free deep reinforcement-learning algorithms, leveraging recurrent neural network architectures (RNNs), have recently been shown to serve as a strong baseline for a wide range of partially-observable Markov decision processes (POMDPs) (Ni, Eysenbach, and Salakhutdinov 2022). Gradient-based reinforcement-learning (RL) algorithms, such as temporal-difference (TD) methods, have been shown to be sample efficient, and come with formal convergence guarantees when using linear function approximation (Sutton and Barto 2018).
Figure 1: The RTRRL recurrent network architecture implements the Meta-RL framework, as defined in Wang et al. (2018), by feeding the past action $a_{t-1}$ to the CT-RNN alongside the current reward $r_{t}$ and the current observation $o_{t}$. The latent state $h_{t}$ which is computed by the CT-RNN, is then used to produce the next action $a_{t}$ and the next value estimate $\hat{v}_{t}$ through linear mappings.
However, linear functions are not able to infer hidden state variables that are required for solving POMDPs. RNNs, on the other hand, can compensate for the partial observability in POMDPs, since they are capable of aggregating information about the entire sequence of preceding observations.
Neural networks were originally inspired by biological neurons, which are in general recurrently connected, and subject to synaptic plasticity. These long-term changes of synaptic efficacy are mediated by locally accumulated proteins, and a scalar-valued reward signal represented by some neurotransmitter (e.g. dopamine) concentration (Wise 2004).
The ubiquitous backpropagation through time algorithm (BPTT) (Werbos 1990), which is used for training RNNs in practice, appears to be biologically implausible, due to the need of weight transport (Bartunov et al. 2018), and distinct forward and backward phases. Thus, BPTT requires individual neurons to alternate between computation and precisely communicating gradients of their synaptic parameters to each other. With BPTT, RNN-based RL algorithms renounce any claim to biological interpretation.
Biologically plausible methods for computing gradients in RNNs however exist, such as random-feedback online learning (RFLO) (Murray 2019), a simplified versions of real-time recurrent learning (RTRL) (Williams and Zipser 1989).
The main question we asked in this paper was if combining biologically-plausible methods for learning and gradient computation in RNNs, such as TD and RFLO, respectively, would be able to learn and solve POMDP RL tasks?
Taking advantage of the previous work, we are able here to answer the above question in a positive fashion, with a novel approach consisting of three basic building blocks:
@@
    A Meta-RL RNN architecture, whose interface and recurrent connections are such that when trained, they become on their own an actor-critic algorithm,
@@
    A TD($\lambda$) Actor-Critic algorithm exploiting temporal difference and dutch eligibility traces, to train the weights of the Meta-RL network, and
@@
    The Biologically Plausible RFLO online automatic differentiation algorithm, for computing the gradients of the parameters in the Meta-RL network.
@@
We call our new, biologically plausible, online RL approach, real-time recurrent reinforcement learning (RTRRL).
The use of TD($\lambda$) in RTRRL, the most basic version of an actor-critic algorithm with no other enhancements, such as the batched experience replay, is able to solve discrete as well as continuous POMDP tasks, in a biologically-plausible fashion. Our experimental results show that, when replacing the RFLO optimization of RTRRL with the BPTT optimization, one achieves a similar accuracy, but in many cases at a much slower convergence time. Moreover, replacing RFLO with RTRL optimization considerably increases time complexity, without significant improvements in accuracy.
Motivation
@@
The main objections for BPTT-plausibility in biological neural networks are the reliance on shared weights, for forward and backward connections, and reciprocal error-transport, a form of propagating back the errors without interfering with neural activity (Bartunov et al. 2018). Another concern is that it requires storing long sequences of exact the exact activation for each cell. (Lillicrap and Santoro 2019)
We slightly rephrase the premises of Bartunov et al. (2018) and Lillicrap and Santoro (2019) in the following fashion. A biologically plausible learning algorithm has to be:
  1. 1.
    Local, up to some low-dimensional reward signal.
  2. 2.
    Online, jointly computing updates and outputs.
  3. 3.
    Without weight transport, in its error computation.
By weight transport, we refer to synapses propagating back error signals mirroring the strength of forward synapses. We will further elaborate on the above requirements for obtaining biological plausibility in Appendix D.
@@
Previous work on feedback alignment (Lillicrap et al. 2014) demonstrated that weight transport is not strictly necessary for training deep neural networks. In particular, they showed that randomly initialized feedback matrices, used for propagating back gradients to previous layers in place of the forward weights, are also leading to acceptable function approximators. Moreover, the forward weights appear to align with the fixed backward weights during training.
@@
Online supervised learning for RNNs was first described by Williams and Zipser (1989), who introduced RTRL as an alternative to backpropagation. More recently, a range of more computationally efficient variants were introduced. (Tallec and Ollivier 2017; Roth, Kanitscheider, and Fiete 2018; Mujika, Meier, and Steger 2018; Murray 2019). Marschall, Cho, and Savin (2020) proposed a unifying framework for all these variants and assessed their biological plausibility while showing that many of them can perform on-par with BPTT. RFLO (Murray 2019), stands out due to its biological update rule, that still performed well in the experiments of Marschall, Cho, and Savin (2020).
Another motivation for investigating online training algorithms of neural networks is the promise of energy-efficient neuromorphic hardware(Zenke and Neftci 2021). The aim is to produce bionic integrated circuits, mimicking biological neurons. However, they also require biologically plausible update rules to enable efficient training.
@@
Our work shows that online reinforcement learning with RNNs is possible with RTRRL-RFLO, which fulfills all our premises for biologically plausible learning. Our algorithm combines online RNN gradient computation with TD($\lambda$). We create a fully online learning neural controller that does not require multi-step unrolling, weight transport or communicating gradients over long distances. Our algorithm succeeds in learning a policy from one continuing stream of experiences alone, meaning that no replay buffer is used.
To the best of our knowledge, the proposed RTRRL, is the first fully biologically plausible reinforcement learning algorithm for rate-based recurrent neural networks.
@@
In this section we present RTRRL, our novel real-time recurrent RL approach to POMDPs. As mentioned, this approach consists of: 1) A Meta-RL actor-critic RNN architecture, which after training becomes on its own an actor-critic algorithm, 2) A TD($\lambda$)-Actor-Critic learning algorithm, exploiting temporal-difference and dutch eligibility traces, to train the weights of the Meta-RL network, and 3) A biologically-plausible real-time RFLO optimization technique, computing the gradients of the parameters in the MetaRL network.
@@
In the following subsections, we provide a gentle and self-contained introduction to each of the constituent parts of our RTRRL framework. We then put all pieces together, and discuss the pseudocode of the overall RTRRL approach.
The Meta-RL RNN Architecture
The Meta-RL actor-critic RNN architecture used by RTRRL is shown in Figure 1. It features a RNN with linear output layers for the actor and critic functions. At each step, the RNN computes an estimated latent state and the two linear layers compute the next action and the next value, from the latent state, respectively. Since the synaptic weights of the network are trained (slowly) to choose the actions with most value, and the network states are also updated during computation (fast) towards the same goal, this architecture is also called a Meta-RL architecture (Wang et al. 2018).
As shown by Wang et al. (2018), a Meta-RL RNN can be trained to solve a general family of parameterized tasks, where all instances follow a common structure, and that the activations in trained RNNs mimic dopaminergic reward prediction errors (RPEs) measured in primates. They also show that the network is capable of inferring the underlying parameters of each new task and subsequently solve unseen instances after training. In particular, they use LSTM cells (Hochreiter and Schmidhuber 1997), receiving the observation, past action, and reward as inputs, and computing a latent state from which the next action and value estimate are inferred by a linear mapping. In Figure 1 we use the architecture of that work, but replace the LSTM with a continuous-time recurrent neural network (CT-RNN).
Continuous-Time RNN
@@
(Funahashi and Nakamura 1993) This type of RNN can be interpreted as rate-based model of biological neurons. They have the following state dynamics:
$\dot{h}=\tau^{-1}\left(-h+\varphi(W_{I}x+W_{R}h+b)\right)$
where factor $\tau$ is the time constant, $\varphi$, a non-linear activation function, $x$, the input, $W_{I},W_{R}$, the synaptic weights for input and states, respectively, and $b$ is the bias.
Wang et al. (2018) argue that, the Meta-RL RNN effectively learns to implement an inner RL algorithm, that empowers it to learn about, and solve tasks on its own. Behrens et al. (2018) argue that, the timescale of plasticity expressed through dopaminergic synapses in the prefrontal cortex, is magnitudes below the time needed for primates to solve tasks, hinting at a hierarchically layered RL algorithms reminiscent of the framework of Wang et al. (2018).
The Actor-Critic Learning Algorithm
As a learning algorithm, RTRRL also uses an Actor-Critic architecture with eligibility traces. Here, the actor and the critic (sharing the common recurrent CT-RNN backbone), are used to compute the synaptic-weights updates, ensuring that for the current observation and reward, and previous action, the network outputs the most valuable action.
In particular, the value produced by the critic, is used to computing the TD error and the forward dutch eligibility traces, which are thereafter used to properly update the synaptic weights. Hence, the RL algorithm used by RTRRL ia an Actor-Critic-TD($\lambda$) algorithm. In order to understand it, we first discuss TD learning, and then eligibility traces.
@@
TD Learning relies only on local information by bootstrapping (Sutton and Barto 2018). It is on-line, which makes it applicable to a wider range of problems, as it does not rely on completing an entire episode prior to computing the updates. After each action, the reward $r_{t}$, and past and current experience $s_{t}$ and $s_{t+1}$ are used to compute the TD error $\delta$ of value estimates $\hat{v}_{\theta}(s)$. The following Equation defines the TD-error:
@@
The value function is learned by regression towards the bootstrapped target. Updates can be computed by taking the gradient of the value approximator $\theta_{t+1}\leftarrow\theta_{t}+\eta\delta_{t}\nabla_{\theta}\hat{v}_{\theta_{t}}(s_{t})$, where $\eta$ is a small step size. In order to also learn behavior, we use an Actor-Critic (AC) policy gradient method. In AC algorithms, the actor computes the actions, and the critic evaluates the actor’s decisions. In particular, the value $\hat{v}_{\theta_{t}}(s_{t})$ of the current state $s_{t}$ is compared to its value $r\,{+}\,\gamma\hat{v}_{\theta_{t}}(s_{t+1})$ after taking an action $a_{t}$, that is, after improving its knowledge by getting a reward $r$ and moving to state $s_{t+1}$. This leads to the the TD-error discussed above.
The policy (actor) is in this case a parameterized function $\pi_{\varphi}$ that maps state $s$ to a distribution of actions $p(a)$. Parameters $\varphi$ are in most cases the weights of a neural network. The policy $\pi_{\varphi}$ is trained using gradient ascent. By taking the gradient of the policy with respect to the log action probability, and multiplying with the TD-error, we can compute the updates $\varphi\leftarrow\varphi+\alpha\delta\nabla_{\varphi}\log\pi_{\varphi}(a)$. Intuitively, this tells the actor to increase the probability for the chosen action whenever the RPE is positive, that is, when the reward was better than predicted. Likewise, when the reward was worse than predicted, the action probability is lowered.
The TD-error is a measure for the accuracy of the reward-prediction, acting as RPE. Given its importance, it is used to update both the actor and the critic, by acting as a reinforcement signal (Sutton and Barto 2018).
@@
The algorithm just described is known as TD(0). It is impractical when dealing with delayed rewards in an online setting, since value estimates need to be updated backwards in order to factor in temporal dependencies. ETs are a way of factoring-in future rewards. The idea is to keep a running average of each synapse’s contribution to the network’s output. This can be thought of as a short-term memory, paralleling the long-term one represented by the synapse weights. ETs unify and generalize Monte-Carlo and TD methods (Sutton and Barto 2018).
Particularly, TD($\lambda$) makes use of ETs. Weight updates are computed by multiplying the trace with the TD-error $\delta$. By accumulating the gradients of the state-value function, each weight’s contribution is recorded. The trace $e^{\theta}$ decays with factor $\gamma\lambda$ where $\gamma$ is the discount factor:
$e_{t}^{\theta}$ $=\gamma\lambda\ e_{t-1}^{\theta}+\nabla_{\theta}\hat{v}_{\theta_{t}}(s_{t})$ (2)
${\theta_{t+1}}$ $\leftarrow{\theta_{t}}+\eta^{\theta}\ \delta_{t}\ e_{t}^{\theta}$
Since we use a parameterized linear-function approximation $\hat{v}_{\theta}(s_{t})\,{=}\,w^{\top}s_{t}$, with parameters $w$ as in the original TD($\lambda$), the gradient of the loss with respect to $w$, is simply $\nabla_{w}\hat{v}_{\theta}\,{=}\,s_{t}$. A later refinement of linear TD($\lambda$) called True online TD($\lambda$) (Seijen and Sutton 2014; Sutton and Barto 2018), improves the accuracy of updates done to the value function. The resulting trace is called a ”dutch” trace. Here, calculating the TD-error can be done without using the same set of weights twice. Hyperparameter $\alpha$ couples the rate of change of the ET with the optimization step size.
$\delta_{t}$ $=r_{t}+\gamma\hat{v}_{\theta_{t}}(s_{t+1})-\hat{v}_{\theta_{t-1}}(s_{t})$ (3)
$e^{\theta}_{t}$ $=\gamma\lambda e^{\theta}_{t-1}+\alpha s_{t-1}-\alpha\gamma\lambda\left[e^{\theta\top}_{t-1}s_{t-1}\right]s_{t-1}$
$\theta_{t+1}$ $=\theta_{t}+\delta_{t}e^{\theta}_{t}+\alpha\left[\theta_{t-1}^{\top}s_{t-1}-\theta_{t}^{\top}s_{t-1}\right]s_{t-1}$
Linear TD methods come with provable convergence guarantees (Sutton and Barto 2018). However, the simplicity of the function approximator fails at accurately representing complex functions needed for solving harder tasks. Replacing the linear functions with non-linear DNNs can introduce inaccuracies in the optimization. However in practice, multi-layer perceptrons (MLPs) can lead to satisfactory results, e.g. on the Atari benchmarks (Daley and Amato 2020).
The Biological-Optimization Algorithm
The optimization algorithm in RTRRL computes the gradients of the synaptic weights, in a biologically plausible fashion. This has to be done for the Meta-RL RNN, in an efficient online fashion as discussed in Section Motivation. For this purpose, RTRRL is using the RFLO algorithm. However, to asses the loss in precision of RTRRL while using RFLO, we also plug RTRL and BPTT in our algorithm, instead of RFLO. In Figure 2, we show how the gradients are passed back to each state of the Meta-RL RNN for RFLO.
@@
RTRL was proposed as an RNN online-optimization algorithm for infinite horizons (Williams and Zipser 1989). This is made possible by estimating the gradient of network parameters in each step, and computing an estimate of the error-vector during the feedforward computation. As in stochastic gradient descent, using a small learning rate $\eta$ can mitigate the noise introduced by estimating the true gradient. One big advantage of RTRL-based algorithms, is that the computation time of an update-step is constant in the number of task steps. Nonetheless, RTRL has a much worse computation complexity than BPTT, and is therefore not used in practice.
The update rule used in RTRL is derived as follows. Given a dataset consisting of the multivariate time-series $x_{t}\,{\in}\,\mathbb{R}^{I}$ of inputs and $y_{t}\,{\in}\,\mathbb{R}^{O}$ of labels, we want to minimize some loss function $\mathcal{L}_{\theta}=\sum_{t=0}^{T}L_{\theta}(x_{t},y_{t})$ by gradient descent. This is achieved by taking small steps in the direction of the negative gradient of the total loss:
$\Delta\theta=-\eta\nabla_{\theta}\mathcal{L}_{\theta}=-\eta\sum_{t=0}^{T}\nabla_{\theta}L_{\theta}(t)$ (4)
We can compute the gradient of the loss at each individual step as $\Delta\theta(t)=\nabla_{\theta}L_{\theta}(t)=\nabla_{\theta}\hat{y}_{t}\nabla_{\hat{y}_{t}}L_{\theta}(t)$ with $\hat{y}_{t}$ being the output of the RNN at timestep $t$. A common choice in supervised learning is the MSE $L_{\theta}(t)=\frac{1}{2}\langle\varepsilon_{t},\varepsilon_{t}\rangle$ where $\varepsilon_{t}=\hat{y}(t)-y(t)\in\mathbb{R}^{O}$ and $\nabla_{\hat{y}_{t}}L_{\theta}(t)=\varepsilon_{t}$. When employing an RNN with linear output mapping the gradient of the model output can be further expanded into $\nabla_{\theta_{R}}\hat{y}_{t}=\nabla_{\theta_{R}}h_{t}\nabla_{h_{z}}\hat{y}_{t}$. RTRL is defined for the special class of Neural ODEs discussed in the previous section.
We calculate the gradient of the Neural ODE unit’s activation with respect to the parameters $\theta$, recursively. To this end, we define the immediate Jacobian $\bar{J}_{t}\,{=}\,\nabla_{\theta}f(x_{t},h_{t})$, with $\nabla_{\theta}$ being the partial derivative with respect to $\theta$, and the approximate Jacobian trace $\hat{J}_{t}$, We will denote the total derivative of $f$ wrt. $\theta$ as $\frac{d}{d\theta}f(x_{t},h_{t})$.
Figure 2: Schematics showing how gradients are passed back to each state of the RNN. The network consists of Linear TD($\lambda$) and a shared RNN body. Gradients with respect to actor and critic losses are propagated back towards $h_{t}$ and $h_{t-1}$ respectively.
@@
$=\frac{d}{d\theta}h_{t}+\frac{d}{d\theta}f(x_{t},h_{t})$
$=\hat{J}_{t}+\nabla_{\theta}f(x_{t},h_{t})+\frac{d}{d\theta}h_{t}\nabla_{h_{t}}f(x_{t},h_{t})$
$=\hat{J}_{t}+\bar{J}_{t}+\hat{J}_{t}\nabla_{h_{t}}f(x_{t},h_{t})$

### v3

@@ Abstract
We introduce a biologically plausible RL framework for solving tasks in partially observable Markov decision processes (POMDPs). The proposed algorithm combines three integral parts: (1) A Meta-RL architecture, resembling the mammalian basal ganglia; (2) A biologically plausible reinforcement learning algorithm, exploiting temporal difference learning and eligibility traces to train the policy and the value-function; (3) An online automatic differentiation algorithm for computing the gradients with respect to parameters of a shared recurrent network backbone. Our experimental results show that the method is capable of solving a diverse set of partially observable reinforcement learning tasks. The algorithm we call real-time recurrent reinforcement learning (RTRRL) serves as a model of learning in biological neural networks, mimicking reward pathways in the basal ganglia.
Codehttps://github.com/FranzKnut/RTRRL-AAAI25
@@ Introduction
Artificial neural networks were originally inspired by biological neurons, which are in general recurrently connected, and subject to synaptic plasticity. These long-term changes of synaptic efficacy are mediated by locally accumulated proteins, and a scalar-valued reward signal represented by neurotransmitter concentrations (e.g. dopamine) (Wise 2004). The ubiquitous backpropagation through time algorithm (BPTT) (Werbos 1990), which is used for training RNNs in practice, appears to be biologically implausible, due to distinct forward and backward phases and the need for weight transport (Bartunov et al. 2018). With BPTT, RNN-based algorithms renounce their claim to biological interpretation. Biologically plausible methods for computing gradients in RNNs do however exist. One algorithm of particular interest is random feedback local online learning (RFLO) (Murray 2019), an approximate version of real-time recurrent learning (RTRL) (Williams and Zipser 1989). Similarly, Linear Recurrent Units (LRUs) (Zucchet et al. 2023) allow for efficient computation of RTRL updates.
Gradient-based reinforcement-learning (RL) algorithms, such as temporal-difference (TD) methods, have been shown to be sample efficient, and come with formal convergence guarantees when using linear function approximation (Sutton and Barto 2018). However, linear functions are not able to infer hidden state variables that are required for solving POMDPs. RNNs, can compensate for the partial observability in POMDPs by aggregating information about the sequence of observations. Model-free deep reinforcement learning algorithms, leveraging recurrent neural network architectures (RNNs), serve as strong baselines for a wide range of partially-observable Markov decision processes (POMDPs) (Ni, Eysenbach, and Salakhutdinov 2022). Contemporary RL algorithms further renounce biological plausibility due to the fact that updates are computed after collecting full trajectories, when future rewards are known.
The question we asked in this paper was whether using a biologically plausible method for computing the gradients in RNNs, such as RFLO, in conjunction with a biologically plausible online RL method, such as TD($\lambda$), would be able to solve partially observable reinforcement learning tasks.
Taking advantage of previous work, we are able to answer the above question in a positive fashion. In summary, our proposed approach consists of three basic building blocks:
Figure 1: RTRRL uses a Meta-RL RNN-backbone which receives observation $o_{t}$, previous action $a_{t-1}$ and reward $r_{t}$, computing the latent vector $h_{t}$ from which the action $a_{t}$ and the value estimate $\hat{v}_{t}$ are computed via linear functions.
@@
    A Meta-RL RNN architecture, resembling the mammalian basal ganglia depicted in figure 1,
@@
    The TD($\lambda$) RL algorithm, exploiting backwards-oriented eligibility traces to train the weights of the RNN.
@@
    Biologically-plausible RFLO or diagonal RTRL, for computing the gradients of RNN-parameters online.
@@
We call our novel, biologically plausible RL approach real-time recurrent reinforcement learning (RTRRL).
@@
In the Appendix, we argue in depth why BPTT is biologically implausible. Here, we summarize the two main objections: First, BPTT’s reliance on shared weights between the forward and the backward synapses; Second, reciprocal error-transport, requiring propagating back the errors without interfering with the neural activity (Bartunov et al. 2018); Third, the need for storing long sequences of the exact activation for each cell (Lillicrap and Santoro 2019).
@@
Previous work on Feedback Alignment (FA) (Lillicrap et al. 2016) demonstrated that weight transport is not strictly required for training deep neural networks. Particularly, they showed that randomly initialized feedback matrices, used for propagating back gradients to previous layers in place of the forward weights, suffice for training acceptable function approximators. They showed that the forward weights appear to align with the fixed backward weights during training.
@@
Online training of RNNs was first described by Williams and Zipser (1989), who introduced real-time recurrent learning (RTRL) as an alternative to BPTT. More recently, a range of computationally more efficient variants were introduced (Tallec and Ollivier 2018; Roth, Kanitscheider, and Fiete 2018; Mujika, Meier, and Steger 2018; Murray 2019). Marschall, Cho, and Savin (2020) proposed a unifying framework for all these variants and assessed their biological plausibility while showing that many of them can perform on-par with BPTT. RFLO (Murray 2019), stands out due to its biologically plausible update rule. LRUs (Orvieto et al. 2023) have gained much attention lately, as they were shown to outperform Transformers on tasks that require long-range memory. Importantly, their diagonal connectivity reduces the complexity of RTRL trace updates, enabling them to compete with BPTT. (Zucchet et al. 2023).
Our second objection to biological plausibility of state-of-the-art RL algorithms is the use of multi-step returns in Monte-Carlo methods. Aggregating reward information over multiple steps helps reducing bias of the update target for value learning. However, in a biological agent, this requires knowledge of the future. While gathering some dust, TD($\lambda$) is fully biologically plausible due to the use of a temporal-difference target and backwards-oriented eligibility trace (Sutton and Barto 2018). Thus, the biological plausibility of RTRRL relies on three main building blocks: (1) The basal-ganglia inspired Meta-RL RNN architecture, (2) The pure backward-view implementation of TD($\lambda$), and (3) The RFLO automatic-differentiation algorithm or LRU RNNs trained with RTRL, as an alternative to BPTT.
@@
This work demonstrates that online reinforcement learning with RNNs is possible with RTRRL, which fulfills all our premises for biologically plausible learning. We create a fully online learning neural controller that does not require multi-step unrolling, weight transport or communicating gradients over long distances. Our algorithm succeeds in learning a policy from one continuing stream of experience alone, meaning that no batched experience replay is needed. Our experimental results show that the same architecture, when trained using BPTT, achieves a similar accuracy, but entailing a worse time complexity. Finally, we show that the use of FA does not diminish performance in many cases.
@@
In this section, we provide a gentle and self-contained introduction to each of the constituent parts of the RTRRL framework, namely the RNN models used, the online TD($\lambda$) actor-critic reinforcement learning algorithm, and RTRL as well as RFLO as a biologically plausible method for computing gradients in RNNs. We then put all pieces together and discuss the pseudocode of the overall RTRRL approach.
Continuous-Time RNN.
@@
Introduced by Funahashi and Nakamura (1993), this type of RNN can be interpreted as rate-based model of biological neurons. In its condensed form, a CT-RNN with $N$ hidden units, $I$ inputs, and $O$ outputs has the following latent-state dynamics:
[…]
Linear Recurrent Units (LRUs).
As a special case of State-Space Models (Gu, Goel, and Re 2021), the latent state of this simple RNN model is described by a linear system:
$\dot{h}_{t}=Ah_{t}+Bx_{t}\qquad\quad y_{t}=\Re\left[Ch_{t}\right]+Dx_{t}$ (2)
where $A$ is a diagonal matrix $\in\mathbb{C}^{N\times N}$ and $B,C,D$ are matrices $\in\mathbb{C}^{N\times I},\mathbb{C}^{O\times N}$ and $\mathbb{R}^{O\times I}$ respectively. Note that the hidden state $h_{t}$ is a complex-valued vector $\in\mathbb{C}^{N}$ here. For computing the output $y_{t}$, the real part of the hidden state is added to the input $x_{t}$ at time $t$. The name Linear Recurrent Unit was introduced in the seminal work of Orvieto et al. (2023). LRUs gained a lot of attention recently as they were shown to perform well in challenging tasks. The linear recurrence means that updates can be computed very efficiently.
The Meta-RL RNN Architecture.
The actor-critic RNN architecture used by RTRRL is shown in figure 1. It features a RNN with linear output layers for the actor and critic functions. At each step, the RNN computes an estimated latent state and the two linear layers compute the next action and the value estimate, from the latent state, respectively. Since the synaptic weights of the network are trained (slowly) to choose the actions with most value, and the network states are also updated during computation (fast) towards the same goal, this architecture is also called a Meta-RL. As shown by Wang et al. (2018), a Meta-RL RNN can be trained to solve a family of parameterized tasks where instances follow a hidden structure. They showed that the architecture is capable of inferring the underlying parameters of each task and subsequently solve unseen instances after training. Furthermore, they showed that the activations in trained RNNs mimic dopaminergic reward prediction errors (RPEs) measured in primates. RTRRL replaces the LSTMs used in Wang et al. (2018) with CT-RNNs, allowing the use of RFLO as a biologically plausible method for computing the gradients of the network’s parameters, or with LRUs which allow for efficient application of RTRL.
@@
TD Learning is a RL method that relies only on local information by bootstrapping (Sutton and Barto 2018). It is online, which makes it applicable to a wide range of problems, as it does not rely on completing an entire episode prior to computing the updates. After each action, the reward $r_{t}$, and past and current states $s_{t}$ and $s_{t+1}$ are used to compute the TD-error $\delta$:
@@
where $\hat{v}_{\theta}(s)$ are value estimates. The value-function is learned by regression towards the bootstrapped target. Accordingly, updates are computed by taking the gradient of the value-function and multiplying with the TD-error $\delta_{t}$: $\theta_{t+1}\leftarrow\theta_{t}+\eta\delta_{t}\nabla_{\theta}\hat{v}_{\theta_{t}}(s_{t})$, where $\eta$ is a small step size.
Policy Gradient.
In order to also learn behavior, we use an actor-critic (AC) policy gradient method. In AC algorithms, the actor computes the actions, and the critic evaluates the actor’s decisions. The actor (policy) is in this case a parameterized function $\pi_{\varphi}$ that maps state $s$ to a distribution of actions $p(a|s)$. The policy $\pi_{\varphi}$ is trained using gradient ascent, taking small steps in the direction of the gradient, with respect to the log action probability, multiplied with the TD-error. Particularly, the updates take the form $\varphi\leftarrow\varphi+\alpha\delta\nabla_{\varphi}\log\pi_{\varphi}(a)$. Intuitively, this aims at increasing the probability for the chosen action whenever $\delta_{t}$ is positive, that is, when the reward was better than predicted. Conversely, when the reward was worse than expected, the action probability is lowered. The TD-error is a measure for the accuracy of the reward-prediction, acting as RPE. Given its importance, it is used to update both the actor and the critic, acting as a reinforcement signal (Sutton and Barto 2018). Note the difference between the reward-signal $r$ and the reinforcement-signal $\delta$: if the reward $r$ is predicted perfectly by $\hat{v}_{\theta}$, no reinforcement $\delta$ takes place, whereas the absence of a predicted reward $r$ leads to a negative reinforcement $\delta$. Although AC was devised through mathematical considerations, the algorithm resembles the dopaminergic feedback pathways found in the mammalian brain.
Figure 2: Schematics showing how gradients are passed back to the RNN (yellow). Gradients of the actor (red) and critic (green) losses are propagated back towards $h_{t}$ and $h_{t-1}$ respectively.
@@
The algorithm just described is known as TD(0). It is impractical when dealing with delayed rewards in an online setting, since value estimates need to be propagated backwards in order to account for temporal dependencies. ETs are a way of factoring in future rewards. The idea is to keep a running average of each parameter’s contribution to the network output. This can be thought of as a short-term memory, paralleling the long-term one represented by the parameters. ETs unify and generalize Monte-Carlo and TD methods (Sutton and Barto 2018). Particularly, TD($\lambda$) makes use of ETs. Weight updates are computed by multiplying the TD-error $\delta$ with the trace that accumulates the gradients of the state-value-function. The trace $e^{\theta}$ decays with factor $\gamma\lambda$ where $\gamma$ is the discount factor:
$e_{t}^{\theta}=\gamma\lambda\ e_{t-1}^{\theta}+\nabla_{\theta}\hat{v}_{\theta_{t}}(s_{t})\qquad\qquad{\theta}\leftarrow{\theta}+\eta^{\theta}\ \delta_{t}\ e_{t}^{\theta}$ (4)
Since in RTRRL we use a linear value-function $\hat{v}_{\theta}(s_{t})\,{=}\,w^{\top}s_{t}$ with parameters $w$, like in the original TD($\lambda$), the gradient of the loss with respect to $w$ is simply $\nabla_{w}\hat{v}_{\theta}\,{=}\,s_{t}$. Linear TD($\lambda$) comes with provable convergence guarantees (Sutton and Barto 2018). However, the simplicity of the function approximator fails to accurately fit non-linear functions needed for solving harder tasks. Replacing the linear functions with neural networks will introduce inaccuracies in the optimization. However in practice, multi-layer perceptrons (MLPs) can lead to satisfactory results, for example on the Atari benchmarks (Daley and Amato 2019). The gradients of the synaptic weights in the shared RNN are computed in an efficient, biologically plausible, online fashion as discussed in the Introduction. To this end, RTRRL is using LRUs trained with RTRL, or CT-RNNs trained with RFLO, an approximation of RTRL. The gradients of the actor and the critic with respect to the RNN’s hidden state are combined and propagated back using Feedback Alignment. In figure 2, we show how the gradients are passed back to the RNN for RFLO.
@@
RTRL was proposed by Williams and Zipser (1989) as an RNN online optimization algorithm for infinite horizons. The idea is to estimate the gradient of network parameters during the feedforward computation, and using an approximate of the error-vector to update the weights in each step. Bias introduced due to computing the gradient with respect to the dynamically changing parameters is mitigated by using a sufficiently small learning rate $\eta$. The update rule used in RTRL is introduced shortly.
Given a dataset consisting of the multivariate time-series $x_{t}\,{\in}\,\mathbb{R}^{I}$ of inputs and $y_{t}\,{\in}\,\mathbb{R}^{O}$ of labels, we want to minimize some loss function $\mathcal{L}_{\theta}=\sum_{t=0}^{T}L_{\theta}(x_{t},y_{t})$ by gradient descent. This is achieved by taking small steps in the direction of the negative gradient of the total loss:
$\Delta\theta=-\eta\nabla_{\theta}\mathcal{L}_{\theta}=-\eta\sum\nolimits_{t=0}^{T}\nabla_{\theta}L_{\theta}(t)$ (5)
We can compute the gradient of the loss as $\Delta\theta(t)=\nabla_{\theta}L_{\theta}(t)=\nabla_{\theta}\hat{y}_{t}\nabla_{\hat{y}_{t}}L_{\theta}(t)$ with $\hat{y}_{t}$ being the output of the RNN at timestep $t$. When employing an RNN with linear output mapping, the gradient of the model output can be further expanded into $\nabla_{\theta_{R}}\hat{y}_{t}=\nabla_{\theta_{R}}h_{t}\nabla_{h_{z}}\hat{y}_{t}$. The gradient of the RNN’s state with respect to the parameters $\theta$ is computed recursively. To this end, we define the immediate Jacobian $\bar{J}_{t}\,{=}\,\nabla_{\theta}f(x_{t},h_{t})$, with $\nabla_{\theta}$ being the partial derivative with respect to $\theta$. Likewise, we introduce the approximate Jacobian trace $\hat{J}_{t}\approx\frac{d}{d\theta}f(x_{t},h_{t})$, where $\frac{d}{d\theta}$ is the total derivative and $f(x_{t},h_{t})=\dot{h}_{t}$ for the respective RNN model used.